# CriticBench: Evaluating Large Language Models as Critic

### v3

@@ Appendix G Details in Objective Evaluation
Appendix H Case Study
This section shows some cases of subjective evaluation of feedback, comparison-based feedback, and correction critique dimensions.
H.1 Case Study Proves Reference Critique
# Question: I like to host guests at my home from time to time, and I am gathering recipes of different dishes and drinks to keep things interesting. I am interested in trying some Danish dishes. Can you give me a recipe for Koldskål? # Generated Response to be Evaluated: Here is a recipe for Koldskål: Ingredients: 4 Quart Jar, 3/4 cup chopped onion 1/2 cup chopped carrot 1/2 cup chopped celery 1 1/2 cups chopped cabbage 1/2 cup chopped mushrooms 1/2 cup chopped tomatoes 1/2 cup chopped zucchini 1/2 cup chopped … (multiple dull repetitions) # Reference Feedback: The response provided is not satisfactory and contains several issues. 1. Factual Error: The recipe provided is not for Koldskål, a traditional Danish dish. Koldskål is a sweet dish typically made with buttermilk, eggs, sugar, vanilla, and lemon, often served with small biscuits called kammerjunkere. The ingredients listed in the response do not reflect this. 2. Relevance: The response does not answer the user’s request for a Koldskål recipe. Instead, it provides a list of ingredients that appear to be for a different type of dish entirely. 3. Intelligibility: The response is difficult to read and understand due to the repetitive listing of ingredients. It’s also missing instructions on how to prepare the dish. 4. Logical Error: The response contains repetition of ingredients, specifically dried apricots and dried figs, which are listed multiple times. This is not logical in a recipe context. To improve the response, the model should provide an accurate recipe for Koldskål, including the correct ingredients and preparation instructions. The model should also avoid repetition of ingredients and ensure the response is clear and easy to understand. # Generated Feedback The recipe for Koldskål is not provided. —————————————————————————————————————————————— # Subjective Evaluation w/. Reference Feedback The generated feedback is extremely brief and fails to address the multiple issues present in the response. It only states that the recipe for Koldskål is not provided, which is accurate but not comprehensive. The reference feedback, on the other hand, provides a detailed analysis of the problems, including factual errors, relevance, intelligibility, and logical errors. It also offers constructive suggestions for improvement. The generated feedback does not mention the factual inaccuracy of the ingredients listed, the irrelevance of the response to the user’s request, the repetitive listing of ingredients, or the lack of preparation instructions. It also does not provide any guidance on how the response could be improved. Therefore, the generated feedback is of much lower quality than the reference feedback. Score: 2 —————————————————————————————————————————————— # Subjective Evaluation w/o. Reference Feedback The generated feedback correctly identifies that the response does not provide a recipe for Koldskål. The response lists ingredients that are not related to the traditional Danish dish Koldskål, which is a sweet cold soup typically made with buttermilk, eggs, sugar, vanilla, and lemon. Instead, the response includes a nonsensical list of ingredients that are repeated multiple times and are more in line with a vegetable or fruit salad rather than a Danish dessert soup. The feedback is succinct and directly addresses the main issue with the response, which is the absence of an actual Koldskål recipe. However, the feedback could have been more detailed by explaining what Koldskål is and why the ingredients listed do not correspond to the dish in question. Score: 8
H.2 Subjective Evaluation Cases
The case for feedback subjective evaluation is shown in Figure 7. The generated response contains factual errors: “cool music from 1920 are King of the Jazz Band and In the Mood.”. High-quality reference feedback points out these errors, while Qwen-72B-Chat’s feedback fails to find them, leading to a low-quality score $2$. Figure 8 shows the case for comparison-based feedback subjective evaluation. The case for correction subjective evaluation is shown in Figure 9.
H.3 Objective Evaluation Cases
The case for the feedback dimension is shown in Figure 10, and the case for the comparison-based feedback dimension is shown in Figure 11.
# Evaluation Task You are given: $\star$ a given user query or instruction $\star$ a generated response by a large-scale language model $\star$ a textual feedback analyzes the response $\star$ a reference high-quality feedback as reference # Given Data ### Question: What is some cool music from the 1920s? ### Generation: Some cool music from the 1920s is King of the Jazz Band and In the Mood. ### Reference Feedback (Score of 8): The response contains a factual error. “In the Mood” is not from the 1920s, but was recorded in 1939 by Glenn Miller. Also, “King of the Jazz Band” is not a recognizable song title from any era. This could lead to confusion for the user. The model should ensure it provides accurate and recognizable song titles from the 1920s. ### Generated Feedback: The response provides two examples of cool music from the 1920s, which are “King of the Jazz Band” and “In the Mood.” However, it would be more helpful if the response could provide additional context or information about these songs or the genre of music they belong to. For example, mentioning that “King of the Jazz Band” was performed by Paul Whiteman’s orchestra, which was one of the most popular jazz bands of the time, or that “In the Mood” was a hit song by Glenn Miller’s orchestra and became an iconic piece of swing music, would make the response more informative and engaging. Additionally, providing a few more examples of cool music from the 1920s would give the reader a broader understanding of the musical landscape of that era. Overall, while the response is relevant and understandable, it can be improved by adding more depth and context. # Very Important Detail!!! 1. If you think the generated feedback is better then reference feedback, you could provide score higher than 8. 2. please first generate your feedback and explanation for the generated feedback by comparing it with reference feedback, and then you must generate the score following the format ‘Score: x‘, where x is the score range from 1 to 10. ——————————————————————————————————————————————— # Chain-of-thought evaluation by GPT-4: The generated feedback fails to identify the factual inaccuracies present. The reference feedback correctly points out that “In the Mood” is not from the 1920s but from 1939, and it also notes that “King of the Jazz Band” is not a recognizable song title. The generated feedback, on the other hand, does not address these issues at all. Instead, it suggests adding more context and examples to make the response more informative, which, while potentially useful, does not correct the fundamental errors. The generated feedback’s suggestion to provide additional context about the songs and the era is a good one, but it is based on the assumption that the provided song titles are correct and relevant, which they are not. Therefore, the generated feedback is not helpful in improving the translation’s accuracy or fluency and does not align with the high-quality reference feedback. # Score: 2
# Evaluation Task You are given: $\star$ a source sentence in English $\star$ a generated response A by a large-scale language model $\star$ a generated response B by another large-scale language model $\star$ a textual comparison-based feedback analyzes the response $\star$ a reference high-quality comparison-based feedback as reference # Given Data ### Translation: Milhaud also used jazz idioms, as in his Suite for clarinet, violin and piano. ### Translation A: 米哈伊尔德也曾使用爵士乐的音调,如他的长笛、小提琴和钢琴的组曲。 ### Translation B: 米尔豪德也使用爵士乐的习语,如他在长号、小提琴和钢琴的四重奏中所使用的。 ### Reference Feedback (Score of 8): Both translations have inaccuracies. In Translation A, “clarinet” is incorrectly translated as “长笛” (flute). In Translation B, “clarinet” is incorrectly translated as “长号” (trombone). Moreover, in Translation B, “Suite” is incorrectly translated as “四重奏” (quartet). However, Translation A correctly translates “Suite” as “组曲”. Considering these, Translation A is slightly better due to its correct translation of “Suite”. ### Generated Feedback: Translation A is more accurate than Translation B. The original sentence mentions “jazz idioms” which is translated correctly as “爵士乐的音调” in Translation A. However, Translation B uses “爵士乐的习语” which does not convey the same meaning as “idioms.” Additionally, the original sentence mentions “Suite for clarinet, violin and piano” which is translated correctly as “长笛、小提琴和钢琴的组曲” in Translation A, but incorrectly as “长号、小提琴和钢琴的四重奏” in Translation B. Therefore, my preference would be for Translation A as it is more accurate and faithful to the original sentence. # Very Important Detail!!! 1. If you think the generated feedback is better then reference feedback, you could provide score higher than 8. 2. Please first generate your feedback and explanation for the generated feedback by comparing it with reference feedback, and then you must generate the score following the format ‘Score: x‘, where x is the score range from 1 to 10. ——————————————————————————————————————————————— # Chain-of-thought evaluation by GPT-4: The generated feedback correctly identifies that Translation A is more accurate than Translation B. However, it makes a significant factual error by stating that “长笛” is the correct translation for “clarinet,” which is incorrect. The correct translation for “clarinet” should be “单簧管” or “黑管.” This error is critical as it misleads the evaluation of the accuracy of the translations. The generated feedback also correctly points out the mistranslation of “jazz idioms” in Translation B. However, it fails to mention that “爵士乐的音调” in Translation A is also not an entirely accurate translation of “jazz idioms,” as “音调” refers more to musical tones rather than idioms or phrases. Additionally, the generated feedback does not address the error in Translation B where “Suite” is translated as “四重奏” (quartet), which is a significant mistake as it changes the type of musical composition being discussed. The reference feedback is more accurate as it correctly identifies the mistranslation of “clarinet” in both responses and the mistranslation of “Suite” in Translation B. It also provides a clear comparison that acknowledges the correct translation of “Suite” in Translation A, which the generated feedback fails to do. In terms of intelligibility, both feedbacks are clear and easy to understand. They both maintain relevance by addressing the accuracy of the translations in relation to the original sentence. Considering these points, the generated feedback has a factual error and misses a critical translation error, making it less effective than the reference feedback. # Score: 4
Figure 8: One case of evaluating generated comparison-based critique on the translation task. Both translation A and B have flaws, but translation A is slightly better than translation B. The response is high-quality, generated by Qwen-7B-Chat mode. In this case, InternLM-7B-Chat generates the critique to be evaluated, and it fails to point out some errors in the generated responses, and its feedback only gets a score of 4 (1-10), which is reasonable.
# Evaluation Task You are given: $\star$ a source sentence in English $\star$ a generated translation $\star$ a high-quality feedback for generated translation $\star$ a textual comparison-based feedback analyzes the response $\star$ a reference high-quality comparison-based feedback as reference Please carefully analyze the source, translation, and feedback for translation, please generate your correction/revision. # Given Data ### Source: Yet, the first rush of the rebels carried the entrenchments of Speen Hill. ### Translation: 然而,第一批起义者攻占了Speen Hill的防御工事。 ### Reference Correction (Score of 8): 然而,起义者的第一波冲击冲破了斯平山的壕沟。 ### Correction: 然而,第一批起义者攻占了Speen Hill的防御工事。 # Very Important Detail!!! 1. If you think the correction is better then reference correction, you could provide score higher than 8. 2. Please first generate your explanation for the generated correction by comparing it with reference correction, and then you must generate the score following the format ‘Score: x‘, where x is the score range from 1 to 10. ——————————————————————————————————————————————— # Chain-of-thought evaluation by GPT-4: The generated correction “然而,第一批起义者攻占了Speen Hill的防御工事。” addresses the feedback points effectively: 1) Fluency: The generated correction replaces “冲了过去” with “攻占了,” which is more formal and appropriate for the historical context of a military action. This improves the fluency of the sentence. 2) Readability: By using “攻占了” (captured), the generated correction clarifies the action taken by the rebels. It removes the confusion caused by “带着Speen Hill的工事冲了过去” in the original translation, making it clear that the rebels overran the entrenchments rather than physically carrying them. 3) Translationese: The term “带着” is removed in the generated correction, thus eliminating the direct translation issue and making the sentence sound more natural in Chinese. Comparing the generated correction to the reference correction “然而,起义者的第一波冲击冲破了斯平山的壕沟。”, both sentences convey the same meaning and are free from translationese. The reference correction uses “冲击冲破了” (assault broke through), which is also a clear and accurate description of the rebels’ action. The reference correction includes the term “壕沟” (trenches), which is a specific type of entrenchment, while the generated correction uses “防御工事” (defensive fortifications), a more general term. Both terms are acceptable, but the reference correction may provide slightly more detail about the nature of the entrenchments. Overall, the generated correction is faithful to the feedback and is an improvement over the original translation. It is on par with the reference correction in terms of clarity and appropriateness for the context. # Score: 8
# Evaluation Task You are given: $\star$ a source sentence in english $\star$ a generated translation # Given Data ### Source (In English): Milhaud also used jazz idioms, as in his Suite for clarinet, violin and piano. ### Translation: 米尔豪德也使用爵士乐的习语,如他在长号、小提琴和钢琴的四重奏中所使用的。 # Very Important Detail!!! Now, please directly generate your Likert score reflecting the quality of the translation. NOTICE: the generation format should be in ‘Decision: x‘, where x is the Likert score range from 1 to 7, where 1 denotes the very bad quality of translation, and 7 denotes the perfect translation. ——————————————————————————————————————————————— # Score: 4
# Evaluation Task You are given: $\star$ a source sentence in english $\star$ a generated translation A $\star$ a generated translation B # Given Data ### Source (In English): Milhaud also used jazz idioms, as in his Suite for clarinet, violin and piano. ### Translation A: 米哈伊尔德也曾使用爵士乐的音调,如他的长笛、小提琴和钢琴的组曲。 ### Translation B: 米尔豪德也使用爵士乐的习语,如他在长号、小提琴和钢琴的四重奏中所使用的。 # Very Important Detail!!! Please carefully analyze these two translations, and select the one that has better quality by choosing the following options: A: translation A is significantly better than translation B. B: translation B is significantly better than translation A. C: none of the translations are significantly better. NOTICE: directly generate your decision in the format ‘Decision: A/B/C.‘ ——————————————————————————————————————————————— # Generation: Decision: A. Translation A is more accurate and faithful to the original sentence. It correctly translates “jazz idioms” as “爵士乐的音调,” which conveys the intended meaning of Milhaud’s use of jazz elements in his music. On the other hand, translation B mistranslates “Suite for clarinet, violin and piano” as “长号、小提琴和钢琴的四重奏,” which is not an accurate representation of the original work. Additionally, the use of “米尔豪德” instead of “米哈伊尔德” may cause confusion for readers who are familiar with the composer’s name. Overall, translation A is a better quality translation.
Figure 11: One case of comparison-based generation on translation task. The preference label is generated by Qwen-72B-Chat, a very powerful open-source LLM.
Appendix I Complete Results
I.1 Complete Results for Different Qualities
I.2 Complete Results for Different Tasks
[…]

### v4

@@ Appendix I Analysis about Length Bias in Subjective Evaluation
Appendix L Evaluated LLMs
We extensively evaluate widely used open-source and closed-source LLMs of different sizes on CriticEval to understand the current progress in this field, including (1) instruction-tuned LLMs; (2) critique-tuned LLMs; and (3) reward models. To reproduce evaluation results, the greedy search is employed for open-source LLMs, and the temperature factor is set as 0 for closed-source LLMs, i.e., decoding randomness is minimum.
The inference procedures of all these evaluated LLMs in this paper are conducted in an A800 server with 8 GPU cards, each with 80G CUDA memory. The vLLM kwon2023efficient and LMDeploy 2023lmdeploy packages are used to speed up the inference, and the average inference time cost for each LLM is 1.25 hours.
L.1 Instruction-tuned LLMs
For closed-source LLMs, we test GPT-4, Claude, Gemini-Pro, PaLM, GPT-3.5-turbo, etc. For open-source LLMs, we test numerous LLM series including Mistral (jiang2024mixtral, ), LLaMA2 (touvron2023llama, ), Baichuan2 (baichuan2023baichuan2, ), Qwen [^6] Qwen-1.5 serie LLMs are only evaluated in the meta-feedback critique dimension. (qwen, ), InternLM2 (2023internlm, ), WizardLM (xu2024wizardlm, ), Vicuna (vicuna2023, ), Yi, and DeepSeek (deepseek-llm, ), etc.
L.2 Critique-tuned LLMs
Recent works have proven that fine-tuning LLMs on critiques generated by GPT-4 significantly improves LLM’s critique ability (li2024generative, ; jiang2023tigerscore, ; xu-etal-2023-instructscore, ; cui2023ultrafeedback, ). Llama-2-13B fine-tuned on GPT-4’s critique could achieve close critique ability to GPT-4, and even outperform much larger model, like Llama-2-70B. In this paper, we name the LLMs fine-tuned on critiques dataset Critique-tuned LLMs, and we comprehensively evaluate critique-tuned LLMs on our proposed CriticEval: (1) TigerScore (jiang2023tigerscore, ); (2) Auto-J-13B (li2024generative, ); and (3) UltraCM (cui2023ultrafeedback, ).
However, there are two popular critique-tuned LLMs are not evaluated in CriticEval. The reasons are listed as follows: (1) InstructScore (xu-etal-2023-instructscore, ) can only be used to evaluate limited tasks, like data2text and commonsense, thus we donot test InstructScore in our work; (2) Prometheus (kim2024prometheus, ) are not evaluated because of its high dependence on the criteria question, score rubrics and reference answers, which is not fully covered in our benchmark.
L.3 Reward Models
Moreover, we show CriticEval can also be used to evaluate reward models (ouyang2022training, ). There are lots of reward models that can be publicly accseed. We only evaluate three representative reward models, and leave the evaluation on other reward models in our future work (lambert2024rewardbench, ): (1) UltraRM-13B (cui2023ultrafeedback, ); (2) Ziya-7B (fengshenbang, ); (3) SteamSHP (pmlr-v162-ethayarajh22a, ).
Appendix M Interpretable Analysis of the Quality of Textual Critiques
Figure 11: Interpretable analysis of the LLM’s critique ability on subjective evaluation.
Beyond the simplified average scores from 1 to 10 in the subjective evaluation of CriticEval, we also categorize the textual critiques of each LLM into three quality intervals for more interpretable analysis: (1) Low-quality critiques (1-3); (2) Medium-quality critiques (4-6); (3) High-quality critiques (7-10). The results of five representative LLMs on feedback, correction, and comparion-based feedback dimensions are shown in Figure 11. It can be found that GPT-4-turbo exhibits strong critique ability and barely generates low-quality critiques. In contrast, the critiques generated by the Llama-2-70B-Chat are usually low-quality. Besides, the ratio of low-quality critiques generated by some LLMs, like Llama-2-70B-Chat and Qwen-72B-Chat, are very high, indicating that they have a lot of room for improvement.
Appendix N Error Patterns in Responses
In this section, we analyze the details and cases about three kinds of error patterns in responses: (1) obvious error; (2) complex error; (3) subtle error. Specifically, we ask human annotators to summarize the common error cases after they annotate all the textual critiques in CriticEval, and categorize them into obvious error, complex error, and subtle errors in Table 28. It should be noted that these three error patterns may have other specific error cases in nine domains. However, it is difficult to exhaust all the error case. Thus, our motivation is to list, annotate, and analyze as many as possible to ensure the accuracy of our experimental results.
Table 28: Specific error cases in each error pattern and each data domain in our human annotation. Since harmful content is easy to detect, the complex error is very rare in CriticEval.
Domains Obvious Errors Complex Errors Subtle Errors
Translate Spelling mistakes Grammatical errors Clear misuses of words Ambiguous Misalignment with input miss translation Incorrect negative expressions Inappropriate word meaning choices Inaccurate translations of idioms Inappropriate tone expressions Misalignment with the background
Chat Fail to fulfill query Grammatical errors Obvious contradictions Incorrect assumptions Flawed deductions Inconsistent arguments Hallucination Logical reasoning error Oversimplification Slight misunderstandings of query Insensitivity to cultural differences Misalignment of style Missing details
QA Contradiction Obvious Commonsense Error Oversimplification Comprehension Difficulty Logical reasoning error Dependency error Hallucination Subtle Contextual Misunderstanding Missing details required in question Imprecise answer Missing details
Harmlessness Explicit forms of discrimination - Implicit bias
Summary Contradiction Missing key point Fail to fullfill query Contextual misunderstanding Factual error Minor deviations in detail Information integration errors
Math CoT Basic Arithmetic mistakes Data entry errors Formula application errors Logical reasoning errors Misunderstanding of problem conditions Omission of key steps Incorrect assumptions Imprecise approximations Ignoring constraints
Math PoT CodeExec CodeNE Syntax error API/Library usage errors Incorrect input/output format Algorithm or Implementation error Time Limit Exceeded Improper handling of edge cases
Appendix O Likert Score for Responses
Appendix P Visualization of Relationship between Model Scales and Critique Ability
Following previous works (luo2023critique, ), we provide the diagrams to demonstrate the relationship between LLM’s critique ability and model scales, which are shown in Figure 12. It can be easily found that the critique ability of all LLM series steadily increase with the number of the parameters (scale) increasing.
Figure 12: The visualization of scaling law of LLM’s critique ability. Under the same LLM series, the critique ability for LLMs becomes better when the size of the model scale increases.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have well discussed the main claims in Section 4 and Section 6.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are well described and discussed in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experimental results could be easily reproduced because we have already discussed the necessary details in Section 6.1, Appendix H and Appendix K.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We upload dataset and evaluation toolkit of our proposed CriticEval with this submission to reproduce our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have well described all the details about inference procedure in Section 6, Appendix G, Appendix F, Appendix E, Appendix D, Appendix L and Appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the statistical significance of our objective evaluation in CriticEval in Section 5 and Section 6.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The time cost, compute worker, CUDA memory for LLM inference in our proposed CriticEval dataset are well described in Appendix L.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We promise our research conforms with the NeurIPS Code of Ethics, which is listed in Section B, Section D.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We have described the ethical consideration in Appendix B, and our paper doesn’t raise the potential positive or negative societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We have discussed this part of content in Appendix B.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have well discussed the Licenses of existing assets in Appendix D, while we do not discuss more details for the used LLMs in this paper due to the vast number of LLMs.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We upload the documentation and some important information about our new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We have discussed this part of the content carefully in Appendix G.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The participants in our work are asked to evaluate the quality of text generations, which doesn’t raise any potential risks. Therefore, the Institutional Review Board (IRB) approvals are easily obtained for our work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
[…]